OpenStudy (anonymous):

simplify (y^2-1)^2+5y(y^2+2y+1) divided by (y+1)^2 and show the workings please!

jhonyy9 (jhonyy9):

\[((y ^{2}-1)^{2}+5y(y ^{2}+2y+1))/(y+1)^{2}=(((y-1)(y+1))^{2} +5y(y+1)^{2})/(y+1)^{2}=(y+1)^{2}((y-1)^{2)}+5y)/(y+1)^{2}=(y-1)^{2}+5y\]

jhonyy9 (jhonyy9):

\[(y+1)^{2}((y-1)^{2}+5y)/(y+1)^{2}=(y-1)^{2}+5y\]

jhonyy9 (jhonyy9):

hope that you will can understanding all